. Write each polar equation in rectangular coordinates and graph. a. r= 2 cos e b. r = - uri
Answer #1

Solu c2coso we know that, x-r cose r. 2.000 Coro 2 82 iekar. = 1 We know that, 8= r = vX2+y? r? x²ty2 x= x² + y2 20=w²ty2 x²​​​​​​